Given the function:y(x,t)=A1cosk1(x+vt)-A2 sink2(x-vt). Determine whether this a solution to the wave equation


1
Expert's answer
2021-07-19T09:52:20-0400

The wave equation has a form

"\frac{\partial^2y}{\partial t^2}=v^2\frac{\partial^2y}{\partial x^2}"

In our case

"y=A_1\cos k_1(x+vt)-A_2\sin k_2(x-vt)"

We get

"\frac{\partial^2y}{\partial t^2}=-A_1k_1^2v^2\cos k_1(x+vt)\\
+A_2k_2^2v^2\sin k_2(x-vt)""v^2\frac{\partial^2y}{\partial x^2}=v^2\left(-A_1k_1^2\cos k_1(x+vt)\right.\\
\left.+A_2k_2^2\sin k_2(x-vt)\right)"

Hence, the function "y(x,t)" is a solution of the wave equation.